OpenStudy (anonymous):

A recipe that makes 6 servings calls for cup of butter. Nancy modifies the recipe so that it can serve 8. How many cups of butter does Nancy need? A. B. C. D.

OpenStudy (anonymous):

the missing part to question is 1/2 a cup and the answers they give are a. 3/8 b. 2/3 c. 1 1/3 d. 3

OpenStudy (shadowcain):

1/2 ÷ 6 = 1/2 ÷ 6/1 = (1 × 1) / (6 × 2) = 1/12 per serving 1/12 x 8 = 1/12 × 8/1 = (1 × 8) / (12 × 1) = 8/12 = 2/3 To divide fractions you have to flip the second fraction and then multiply. So it will look like this 1/2 * 1/6 and you get 1/12 this is how much you need for 1 serving now multiply that by 8 and you get 8/12 now simplify and you get 2/3 as your answer.